OpenStudy (anonymous):

simplify: 9*9^1/2 (exact answer... radicals as needed)

OpenStudy (dumbcow):

9*3 = 27 ?

jimthompson5910 (jim_thompson5910):

\[\large 9*9^{\frac{1}{2}}=9\sqrt{9}=9*3=27\] So \[\large 9*9^{\frac{1}{2}}=27\]